A 71-year-old man weighing 155 lbs was determined to have a peak VO2 of 45ml/kg/min (1...

  1. A 71-year-old man weighing 155 lbs was determined to have a peak VO2 of 45ml/kg/min (1 pt).

    • What is his oxygen consumption in absolute terms (L/min)?

    • In METs?

    • How many kcals/min is he expending?

  1. 80 kg man did workout and spent 2.5 L/min O2. What intensity exercise did he do (Hint. Convert to METs)? 1pt


  1. Theodore G is 35 years old and normally active. He did rollerblading at 17.5 ml/kg/min

    • What intensity (Calculate METs and specify low, moderate, and high intensity)? 1 pt


  1. A 30-yr-old male participant has a measured maximal HR of 188 bpm and a resting HR of 66 bpm. His training intensity goal is 65%HRR. What is his target heart rate (THR)? 1 pt

  1. A 75kg man completes a maximal Grade exercise testing, and his VO2max is 3.0 L/min. (2 pts)

    • This equals __________ kcal/min.

  • This equals ______ml/kg/min.

  • This equals ______Mets.

  • At what exercise intensities (=target VO2R) does he work in order to be at 60% and 80% VO2R?

  1. A 40-year-old male participant has a measured maximal HR of 175bpm and resting HR of 75 bpm. What is his 60% and 80% of the THR as calculated by the HRR method? (1.0 pts)

    • His resting HR decreases by 10 beats/min after participating in an endurance-training program. Because maximal HR (175 beats/min) is not affected by training, what happens his 60% and 80% of THR (Calculate THR and then explain your answer)?


  1. What is the gross relative oxygen consumption (ml/kg/min) for a 125-lb, 35 yr old female walking on a treadmill at 3.5 mph and a 2% grade? (1 points)



  1. Subject: A male 30 yr of age, 150 lb, Height, 70 inch. VO2max= 60ml/kg/min; desired treadmill grade: 2.5% and Desired Exercise intensity: 75% VO2Reserve. Using Table 6.3 (available on Blackboard), determine running speed (Km/h) on a treadmill (2 point).
Answer #1

Answer:

1):

  • lbs = 0.454 kg
  • So 155lbs = 70.3 Kg.
  • A man have peak VO2 of 45ml/kg/min.

a):

  • It means in 1 kg it has consumption of 45ml/min
  • So in 70.3 kg it has oxygen consumption of
  • 45×70.3 = 3163.5 ml/min.

b):

  • Since 3.5 ml/kg/min = 1MET.
  • So 45 ml/kg/min = 12.85 MET

c):

Using conversion formula

  • MET to kcal/min = (multiply METs × 3.5 ×body weight in kg)/200
  • So 12.85 MET to kcal/min = (12.85 × 3.5 × 70.3)/200
  • =15.80 kcal/min

2):

  • 1 MET(Metabolic Equivalent of Task) = 3.5ml of O2/kg/min is the energy required by the body while sitting quietly and doing nothing.
  • Means 3.5ml of Oxygen is required by the body per kg per min and this Oxygen is the source of energy that makes us sit quietly.
  • If we do some physical activity then more amount of Oxygen will be required to do that task.
  • In this question the person is doing rollerblading at 17.5ml/kg/min. So MET= 17.5/3.5 = 5.
  • Now according to the table as shown below MET ranging from 3-6 are considered as moderate intensity tasks.
  • MET itself is an index of the intensity of the task. Hence, the answer to this question is 5 MET and in moderate range.

3):

Please Rate My Answer...... Thank......u...

NOTE:As per Chegg rules i have to write Actually 1or2 questions only...